• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 08:19:10 on April 17, 2004

    Entry number 118276

    keyword=Beam Time Accounting
    
    	
    Hall A

    04/17/04 0000-0800
    Experiment #: E94-107
    Shift Leader: J. Singh
    Shift Workers: M. Iodice, V. Nelyubin

    00000100020003000400050006000700Total Time
    Acceptable Beam
    in Use
    0.330.330.240.090.790.040.620.683.12
    Beam Available
    but Not in Use
    0.550.480.740.850.070.000.330.253.28
    Beam Not Available
    or Unacceptable
    0.120.190.020.060.140.970.050.071.62
    Accelerator Con-
    figuration Change
    0.000.000.000.000.000.000.000.000.00
    Experiment
    Ready
    0.450.520.250.331.001.000.670.835.05
    Planned Config-
    uration Change
    0.550.480.750.670.000.000.000.002.45
    Unplanned
    Experiment Down
    0.000.000.000.000.000.000.330.170.50
    Hall not
    ready reasons
    Target0.000.000.000.000.000.000.000.000.00
    Data Acq0.000.000.000.000.000.000.250.000.25
    Other0.000.000.000.000.000.000.000.000.00
    HRSL0.000.000.000.000.000.000.080.170.25
    HRSR0.000.000.000.000.000.000.000.000.00
    HALL_OFF0.000.000.000.000.000.000.000.000.00

    Comments: from 0000 to about 0330 Eugene was doing detector studies. he spent some time changing magnet settings and stuff. from 0500 to 0600, we had MCC try to get the OTR working. see 118236. i'll call this BNA b/c we were trying to do this to figure out our energy spread.

    Data Entry by: J. Singh